Strontium,Sr2+, is a heavy metal tht frequently causes concern when it is present in water. Some preliminary calculations are necessary to determine whether it is feasible to precipitate Sr2+ as the sulfate. Given that 10-4 mole/ liter of Sr2+ and 10^-4 mole/liter of SO42- are originally present and that pK = 7.9 FOR SrS04

1) How much Sr2+ will precipitate if 1x10^-3 mole Na2SO4 is added perliter of solution?

2) How much SO42- and Sr2+ remian after precipitation?

3) Assuming that the residulas calculated in part b) are satisfactory, what factors should be studied experimentally to determine the feasibility of the process?

1. Strontium, Sr is a heavy metal that frequently
Answer #1

SrSO4 (s) < -------------> Sr2+ (aq.) + SO42- (aq.)

K = [Sr2+][SO42-]

10-7.9 = (S) (0.0001+0.001)

S = 1.14 * 10-5 M

Initial moles of Sr2+ = 0.0001 mol

Final moles of Sr2+ = 0.0000114 mol

moles of Sr2+ that is precipitated = 0.0001 - 0.0000114 = 0.0000886 mol

(b) Remaining Sr2+ = 0.0000114 mol and SO42- = 0.0011 mol

(c) Because the K of SrSO4 is small it is feasible to precipitate Sr2+ as its sulphate.